R1 L1 10 1H 10 cos(500t-60°) C1 =1uF

The average power at the source IS EQUAL TO WHAT WITH DEMONSTRATION,?
Answer #1

Islocos 5001-60) 21 RI ML 10 TH Ta IMF Zin = R1 + iWL, I Zin = 10 +5500 - i WC W = 50orad/sec 500x/06 Zin = 10 - 31500,2 Zin